OpenStudy (anonymous):

Convert to degrees 11pi/15 a.)66 b.)132 c.)528 d.)264 would the answer be 528 or 264

OpenStudy (theeric):

How did you get to them?

OpenStudy (theeric):

Just so I know what process you're trying to use.

OpenStudy (luigi0210):

To convert use: 180/pi

OpenStudy (anonymous):

those two answer seem to be right to me because, the others are too low @theEric

OpenStudy (luigi0210):

\[\frac{11 \pi}{5}* \frac{180}{\pi}\]

OpenStudy (anonymous):

its a trap

OpenStudy (anonymous):

don't belive luigi

OpenStudy (luigi0210):

*15

OpenStudy (theeric):

Here's a good way to think about it. A full circle is a full circle. Using a substitution, \(2\pi\) radians is \(360^\circ\). Mathematically,\[2\pi\ [radians]=360^\circ\]\[\qquad\Downarrow\]\[1=\frac{360^\circ}{2\pi\ [rad]}=\frac{180^\circ}{\pi\ [radians]}\] So you can convert like that. I think the 5 is supposed to be a 15 according to the question! Right? That gives you B.

OpenStudy (theeric):

See how I did that, @arria ?

OpenStudy (theeric):

If you ever forget what to do, remember that a circle is a circle. Then you can convert!

OpenStudy (luigi0210):

@Jack174 why are you putting me down mam?

OpenStudy (anonymous):

when you covert do you always use 360 because of the circle @theEric

OpenStudy (luigi0210):

@arria use 180..

OpenStudy (anonymous):

thanks i already have the answer @Luigi0210

OpenStudy (theeric):

\[\frac{11\cancel\pi}{15}\frac{180^\circ}{\cancel\pi\ [rad]}\] Yeah, @arria . The trick is in the converting. See, \(2\pi\ [radians]=360^\circ\) always, right? That's just two different units for measuring angles, where\(2\pi\) radians is a full circle and \(360^\circ\) is a full circle. So \(2\pi\ [radians]=360^\circ\), or \(\pi\ [radians]=180^\circ\) after dividing both sides by \(2\) to make it simple.

OpenStudy (theeric):

So, when you convert, you can use \(\Large \frac{360^\circ}{2\pi\ [rad]}\), or \(\Large \frac{180^\circ}{\pi\ [rad]}\) Just notice you make the units cancel! \[\frac{11\pi}{15}\ \cancel{[rad]}\times\frac{360^\circ}{2\pi\ \cancel{[rad]}}=132^\circ\]

OpenStudy (theeric):

And just remember how I got to the conversion, \(2\pi\ [radians]=360^\circ\). Because it will also get you to \(\Large\frac{2\pi\ [radians]}{360^\circ}\large =1\). And multiplying by \(1\) doesn't change the value, which is why the conversion works. Any questions at all, @arria ?